OpenStudy (anonymous):

The round trip flight took 3.9 h flying time. The plane traveled the 510 mi to the city at 255 mi/h with no wind. How strong was the wind on the return flight? Was the wind a head or tail wind?

OpenStudy (radar):

Let x= speed in mph on the return flight. 510/x=time distace/rate=time Now to figure time for the first half of the flight. distance =510 miles, rate was given as 255 mph. time=510/255 = 2 hours for the first half. Time for the round triip is given as 3.9h. Subtract 2 and you find the time for the return trip. 3.9-2=1.9. Since this is less time for the first part you can answer part of problem now. The wind was a tail wind. Now to the equation 510/x=1.9. 1.9x=510 x=268.421 mph How strong was the wind? 255-268.421=13.42 mph tail wind.

OpenStudy (anonymous):

Thanks!

OpenStudy (radar):

NP, gud luckl wid it.